OpenStudy (anonymous):

exponential form: p^4 x q^4 do i add or multiply the raised numbers?

OpenStudy (anonymous):

well u can do neither because the bases are not the same. if you had q^4xq^4 then u add the raised numbers. but the one base is p and the other is q so u can do neither

OpenStudy (amistre64):

neither; they are unlike bases

OpenStudy (anonymous):

Neither, it's just\[p ^{4}q ^{4}\]

OpenStudy (anonymous):

so the expression cannot be simplified? im just confused when i am supposed to add or multiply these. let me use another example so someone can tell me what the eff im doing wrong.

OpenStudy (amistre64):

think of it as apples and oranges ... you can make a fruit salad out of it, but other than that they remain seperate

OpenStudy (amistre64):

apples^4 + oranges^4 = .... well, thats all it can equal :)

OpenStudy (anonymous):

lol.ok lishfosho

OpenStudy (anonymous):

(8x^3-5x^2+5x)+(3x^3+9x^2-12x) do i multiply or add the exponents?

OpenStudy (anonymous):

the first question can be written as (pq)^4 or p^4q^4

OpenStudy (anonymous):

If you want, you can change it to (PQ)^4 Combining is allowed like that; only if there's no addition involved.

OpenStudy (anonymous):

thats simplified? now im confuzzed.

OpenStudy (anonymous):

second question... open the second bracket and here u add like terms...

OpenStudy (anonymous):

Neither again.

jhonyy9 (jhonyy9):

- so this is (pq)^4

OpenStudy (anonymous):

Often times 'simplify' doesn't actually make things any simpler. They just want you to demonstrate knowledge of some algebraic property and cannot be bothered to tell you which one.

OpenStudy (anonymous):

;p

OpenStudy (anonymous):

For the second question: (8x^3-5x^2+5x)+(3x^3+9x^2-12x) You have to just combine your like terms.

OpenStudy (anonymous):

so when i combine the 8x^3 + 3x^3 would that = 11x^6 or 11x^9 ?

OpenStudy (anonymous):

You cannot do anything with exponents unless you are multiplying or exponentiating.

OpenStudy (anonymous):

Therefore don't change the exponent.

OpenStudy (anonymous):

its just 8x^3 + 3x^3. Leave it as it is.

OpenStudy (anonymous):

Because you are adding, not multiplying.

OpenStudy (anonymous):

11x^3

OpenStudy (anonymous):

right, for the x^3'd terms.

OpenStudy (amistre64):

that was ..... almost rude lol. I gave explanations on how to work it, but i understand your response to it :)

OpenStudy (anonymous):

thank you so much and if i was multiplying, i would multiply the exponents?

OpenStudy (anonymous):

how is that right? i thought you could combine all like terms?

OpenStudy (amistre64):

is 9 the same as 3?

OpenStudy (anonymous):

I am often almost rude. Sorry.

OpenStudy (amistre64):

3^2 is not the same term then as 3^1 x^2 and x are just as unlike termed

OpenStudy (anonymous):

If you are multiplying, you add exponents. \[x^b \cdot x^c = x^{b+c}\] If you exponentiate you multiply exponents \[(x^b)^c = x^{b\cdot c}\]

OpenStudy (anonymous):

Did you finish combining your terms for this one? What'd you get?

OpenStudy (anonymous):

I got 11x^6+4x^4-7x^2

OpenStudy (anonymous):

I am TERRIBLE at math, thank you all for your patience and my problem of NOT understanding.

OpenStudy (anonymous):

No, you don't change the exponents.. remember?

OpenStudy (anonymous):

so 11x^3+4x^2-7x

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Since you're only adding/subtracting the exponents don't change.

OpenStudy (anonymous):

They only change if you multiply, divide, or exponentiate.

OpenStudy (anonymous):

goodness, can you just take my final for me? haha.

OpenStudy (anonymous):

Mathematics is not an innate skill. There's no magic, just hard work and practice. If you do that, you will get better. I promise. =)

OpenStudy (anonymous):

Thank you. So since you help out and don't tell me the answer, I am understanding it much better. With exponential form.(5^4)(5^7)(5^8) would this be 5^19 ?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

YAY. go me! haha Thank you! (I am sure I will be back...haha)